-2X+3Y=14
X+2Y=7 HAVING PROBLEMS SOLVING THIS PLEASE HELP

Answer by elima(1433)

-2X+3Y=14
X+2Y=7
First we will solve the second eq. for x;
x+2y=7
x=7-2y
Now that we solved for x, plug that into eq. 1 and solve for y;
-2(7-2y)+3y=14
-14+4y+3y=14
4y+3y=14+14
7y=28
y=4
Now that we have y, plug tha into eq.2 and find x;
x+2(4)=7
x+8=7
x=-1
So we have;
x=-1
y=4
Hope you understand!
=)